OpenStudy (anonymous):

how come y'' + 4y=0 is homogeneous?

OpenStudy (anonymous):

dy/dx+4y=0 dydx=-4y-------(1) substitutig ky =>y'=-4ky which is not equal to (1) therefore its not a homogeneous equation

OpenStudy (anonymous):

Well go back to the previous section and take a look at Example 7 and Example 8. In those two examples we solved homogeneous (and that’s important!) BVP’s in the form, search for this at http://tutorial.math.lamar.edu/Classes/DE/BVPEvals.aspx

OpenStudy (anonymous):

it says that it is a homogeneous

OpenStudy (anonymous):

isn't because this diff eq has set to zero i.e. g(t)=0

OpenStudy (anonymous):

oopps !!!!! its double differentiated equation y''=-4y dy/y=-4 . dx integrating on both side logy =-4x but still it isnt a homogeneous eq

OpenStudy (anonymous):

A DE is said homogeneous if there is a homogeneous function F such that \[y \prime = F(x,y)\] i.e., I think that you can write this as: y'' = -4y y' = -4yx and F(x,y) = -4yx. Only have to prove that F is homogeneous now.

OpenStudy (experimentx):

I was thinking the same as @bmp but somewhere I remembered, if the function can be defined as F(y/x) then it is called homogeneous differential equation.

OpenStudy (experimentx):

though I am not quite sure about either.

OpenStudy (anonymous):

@experimentX I think this proof suffices, since F(tx, ty) = t*(-4yx) = t*F(x,y), t being a scalar. So F is homogeneous

OpenStudy (anonymous):

I meant, t^2*F(x,y)

OpenStudy (experimentx):

not really sure ... i'm quite illiterate on these stuffs.